OpenStudy (anonymous):

The beginning tennis class has twice as many students as the advanced class. The intermediate class has three more students than the advanced class. How many students are in the advanced class if the total enrollment for the three tennis classes is 39?

OpenStudy (mia99):

b = beginning class a = advanced class i = intermediate class b + a + 1 = 39 b = 2a i = a + 3 now sub 2a in for b and (a + 3) in for i in the first equation... b + a + i = 39 (2a) + a + (a + 3) = 39 -- combine like terms 4a + 3 = 39 4a = 39 - 3 4a = 36 a = 9 ANSWER : C advanced = 9 intermediate = a + 3 = 9 + 3 = 12 beginning = 2a = 2(9)= 18 check... B + a + i = 39 18 + 9 + 12 = 39 39 = 39 (correct)